a truck enters the highway driving 60 mph. a car enters the highway at the same place 8 minutes later and drives 70 mph in the same direction. from the time the car enters the highway, how long will it take the car to pass the truck?
Answer by richwmiller(17219)

Gaining speed 70 -60= 10 mph
Headstart 60*8 =480 miles
Time to catch up 480/10=48 min
Algebraic solution
60(t+8)=70t
60t+480=70t
480=70t-60t
480=10t
48=t
Distance to meeting 48*70=3360 miles